please give me all formulas of states of matter lesson!!!!!!! like all in series like r1/r2=v1/v2 and soo on ???? pls sir

please 
give me all formulas of states of matter lesson!!!!!!!
like all in series like r1/r2=v1/v2 and soo on ????
pls sir 

Grade:11

1 Answers

Preeti
6 years ago
Pv=Nrt is the ideal gas equation.Boyles law - p1v1=p2v2Charlie`s v1/v2=t1/t2Gay lussac p1/p2 t1/t2Avogadro v1/v2=n1/n2Combined gas eq is p1v1/t1=p2v2/t2Partial pressure of gas = mole fraction of gas*p.totalP(dry gas)= p.total(subtract)-aq.tensionPv (mix)=p1v1+p2v2